THE SPEED OF THE PROJECTILE WHEN IT IS AT THE MAXIMUM HEIGHT IS ROOT 2/5 TIMES OF ITS SPEED
WHEN IT IS AT HALF OF THE MAXIMUM HEIGHT THE ANGLE OF PROJECTION IS? 

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
Speed of shot at the most astounding point = ucosα (α is the edge of projection) 
stature of the projectile= u2sin2α/2g 
MAX ht/2= u2sin2α/4g 
Speed at MAX ht/2 = v 
u2-v2=2gh 
=>v2=u2-2gx u2sin2α/4g 
=>v2=u2(1-sin2α/2) 
ucosα=√(2/5) √u^2 (1-sin^2⁡α/2) 
=>cosα=√(2/5) √(1-sin^2⁡α/2) 
=>cos2α=(2-sin2α)/5 
=>5cos2α=1+cos2α 
=>cos2α=1/4 
=>α=60⁰ .
So the prjectijon of angle found is 60 degree.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)